OpenStudy (anonymous):

I need helping working this out really quickly please! Solve: 1/3 + 1/6 = 1/x

OpenStudy (neer2890):

just take l.c.m and find x.

OpenStudy (neer2890):

\[=\frac{2+1 }{ 6 }\]

OpenStudy (anonymous):

I thought the answer was 1/2 but i didn't know if i could just cross multiply

OpenStudy (neer2890):

\[\frac{ 3 }{ 6 }=\frac{ 1 }{ x }\] now cross multiply 3x=6 x=2

OpenStudy (anonymous):

thank you so much